You are given the line "l:\mathbf{r}=(1+2\lambda,-1-\lambda,3\lambda)" , where "\lambda" is a parameter.

Which of the following statements are true?




  1. The point (3,-2,3) lies on the line "l"
  2. The vector (2,−1,3) is parallel to the line.
  3. Line "l" is perpendicular to the line "\mathbf{r}=(2t, 2-t,-1+3t)" , where "t" is a parameter.
  4. The equation of a plane containing line "l" and also passing through the point "(0,2,\u22121)" is "8x+y-5z=7" .
  5. The vector "(2,\u22121,3)" is perpendicular to the line "l" .
1
Expert's answer
2021-10-18T07:14:55-0400

1.If the point (3, -2, 3) lies on the line "l: r=(1+2\lambda,-1-\lambda,3\lambda)", then "(1+2\lambda=3)\land(-1-\lambda=-2)\land(3\lambda =3)"

All three conditions is true when "\lambda= 1", so the point lies on the line "l" when "\lambda = 1"

2.If the vector (2, -1, 3) is parallel to the line, then "\exist t\in R:t*(a, b, c) = (2, -1, 3)", where a, b, c - coefficients before "\lambda" for every coordinate

In our case: a = 2, b = -1, c = 3. So If the vector (2, -1, 3) is parallel to the line, then "\exist t\in R:t*(2, -1, 3) = (2, -1, 3)". We have equility when t = 1, so vector (2, -1, 3) is parallel to the line

3.If two lines are perpendicular, the "a{\scriptscriptstyle 1}*a{\scriptscriptstyle 2}+b{\scriptscriptstyle 1}*b{\scriptscriptstyle 2}+c{\scriptscriptstyle 1}*c{\scriptscriptstyle 2} = 0", where "a{\scriptscriptstyle 1},b{\scriptscriptstyle 1},c{\scriptscriptstyle 1}" and "a{\scriptscriptstyle 2}, b{\scriptscriptstyle 2}, c{\scriptscriptstyle 2}" are coefficients before parameters for every coordinate for first and second line

In our case: "a{\scriptscriptstyle 1}*a{\scriptscriptstyle 2}+b{\scriptscriptstyle 1}*b{\scriptscriptstyle 2}+c{\scriptscriptstyle 1}*c{\scriptscriptstyle 2} = 2*2 + (-1)*(-1)+3*3 = 14 \not = 0\to" lines aren't perpendicular

4.Firstly we find two points that belong to the line l

"M{\scriptscriptstyle 1} = (1, -1, 0)", when "\lambda = 0"

"M{\scriptscriptstyle 2} = (3, -2, 3)", when "\lambda =1"

"M{\scriptscriptstyle 3} = (0, 2, -1)"

Now we have to find the equation of the plane that contains 3 points

det"\begin{pmatrix}
 x-1 & y+1 & z \\
 3-1 & -2+1&3 \\
 0-1&2+1&-1
\end{pmatrix}" = (x-1)*(-2+1)*(-1)+(y+1)*3*(-1)+z*(3-1)*(2+1)-(-1)*(-2+1)*z-(2+1)*3*(x--1)-(-1)*(3-1)(y+1) = x-1-3y-3+6z-z-9x+9+2y+2 = -8x - y +5z +7 = 0

The equation of the plane is "-8x - y +5z +7 = 0 \to 8x+y-5z+7 \to" The statement is true

5.In the second statement we found out that that vector is parallel to the line, so it's not perpendicular


So, the statements 1, 2, 4 are true